OpenStudy (ilovecake):

What is (n-5) times 6/ n divided by 4 when n = 8?

OpenStudy (ilovecake):

\[(n-5)\times6 \over n \div4\]

OpenStudy (haseeb96):

just put n= 8 into the equation u wrote @Ilovecake

OpenStudy (anonymous):

just plug in the numbers and thats it

OpenStudy (ilovecake):

Do I do it like this: 8-5=3 times 6= 18. then on the bottom: 8 divided by 4 is 2 then add 18+2=20? Is 20 the answer?

OpenStudy (anonymous):

you take 6 * n and you get that answer and then you do 6*5 and get that answer then you do n/4 and then you add the top two numbers after you figure out that answer then you divide that and you get the answer I got 39

OpenStudy (ilovecake):

n=8 though

OpenStudy (anonymous):

sorry I did adding instead of subtract sorry

OpenStudy (ilovecake):

oh ok

OpenStudy (haseeb96):



OpenStudy (ilovecake):

the answer is 9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer is 9

OpenStudy (haseeb96):

yes it is correct answer

OpenStudy (ilovecake):

Who do I give the medal to?

OpenStudy (anonymous):

you can give it to @Haseeb96

OpenStudy (ilovecake):

are you sure? okay

OpenStudy (haseeb96):

what ?

OpenStudy (anonymous):

yes im sure

OpenStudy (haseeb96):

i dont need any medal i m here only to help the other

OpenStudy (anonymous):

im also here to help people

OpenStudy (ilovecake):

Can you guys help me with some more

OpenStudy (anonymous):

yea what's the question

OpenStudy (ilovecake):

ok hold on

OpenStudy (ilovecake):

Evaluate. 1 – 3 • 7 + 4 • 2 A. –28 B. –12 C. –20 D. –32

OpenStudy (haseeb96):

write it fromm the start

OpenStudy (ilovecake):

okay.

OpenStudy (ilovecake):

I think it is A or D

OpenStudy (anonymous):

I got A when I did 1 *3+7+4*2

OpenStudy (ilovecake):

ok

OpenStudy (ilovecake):

so the answer is A? for sure?

OpenStudy (anonymous):

yes @Haseeb96 do you think this is right to?

OpenStudy (haseeb96):

please tell me what is the equation ?

OpenStudy (ilovecake):

1 – 3 • 7 + 4 • 2

OpenStudy (haseeb96):

it is not showing to me please write it

OpenStudy (anonymous):

I did 1*3+7+4*2 and I had gotten 28

OpenStudy (ilovecake):

1 - 3 times 7 + 4 times 2

OpenStudy (anonymous):

you get -20 if you do it that way @Ilovecake

OpenStudy (ilovecake):

That is the equation...

OpenStudy (anonymous):

then none of those answers would be correct

OpenStudy (haseeb96):

1-3 =-2 * 7 = -14 4 *2 =8 -14 + 8 -6 is the correct answer

OpenStudy (haseeb96):

can u give me the all the choices ?

OpenStudy (ilovecake):

-6 is not an answer choice

OpenStudy (ilovecake):

A. –28 B. –12 C. –20 D. –32 Refresh the page to see it.

OpenStudy (haseeb96):

please write the all choices and equation also

OpenStudy (ilovecake):

If you see question marks, refresh the page.

OpenStudy (ilovecake):

Evaluate. 1 – 3 • 7 + 4 • 2 A. –28 B. –12 C. –20 D. –32

OpenStudy (anonymous):

see I still think its A because the answers that we have came up with don't match the answers you have besides 28

OpenStudy (ilovecake):

okay. I hope your right. next question..

OpenStudy (anonymous):

ok what's the question

OpenStudy (haseeb96):

B is the correct answer

OpenStudy (anonymous):

how did you get B

OpenStudy (haseeb96):

first multiplication function will be held 1-21 + 8 -20 +8 -12

OpenStudy (haseeb96):

i hope u got it